1. A 2.5-cm-diameter horizontal water jet with a speed of Vj = 40 m/s relative to the ground is deflected by a 60° stationary cone whose bat diameter is 25 cm. Water velocity along the cone varies linearly from zero at the cone surface to the incoming jet speed of 40 m/s at the free surface.



2. A horizontal war jet impinges against a vertical flat plate at 30 ft/s and splashes off the sides in the vertical plane. If a horizontal force of 500 lbf is required to hold the plate against the water stream. determine the volume flow rate of the water.


1
Expert's answer
2021-12-17T18:16:35-0500

"F_J=F_R,"

"P_JA=F_R,"

"\frac 12\rho v^2\cdot \frac{\omega}v=F_R,"

"\frac 12\rho v\omega=F_R,"

"\omega=\frac{2F_R}{\rho v},"

"\omega=0.59~\frac{ft^3}s."